OpenStudy (anonymous):

solve the equation by factoring 24x^2 -30x =0 a. x=54,x=30 b. x=0 c. x=5/4, x=0 d. x=5/4

OpenStudy (anonymous):

Start by taking x as a common factor: \[x(24x-30)=0\] Multiplication of two number is zero, means that either one of them can be zero. So x=0 or \[24x-30=0 \implies x={30 \over 24} \implies x={5 \over 4}\] So the equation has two solutions x=0 and x=5/4

OpenStudy (anonymous):

the other question is different but i think the way u explain this problem helped me solve the last one i just posted thanks

OpenStudy (anonymous):

You're welcome!! And thanks for the medal.